OpenStudy (anonymous):

Last year you mowed grass or shoveled snow for 12 households. You charged $200 a season to mow a yard and $180 a season to shovel snow. You earned $2280 last year. How many households did you mow

OpenStudy (hoblos):

x=number of households mowed y=number of households shoveled snow for x+y=12 200x+180y=2280 from the first equation we get y=12-x replace it in the second 200x + 2160 - 180x = 2280 20x = 120 x=6

OpenStudy (anonymous):

thanks what is y ?

OpenStudy (hoblos):

y=12-x = 6 also it is the number of households shoveled snow for

OpenStudy (anonymous):

Ok thank you !!!!

OpenStudy (hoblos):

my pleasure XD